OpenStudy (anonymous):

factor by grouping a^2+2ab-24b^2

OpenStudy (whpalmer4):

Factor (by grouping) \[a^2 + 2 a b - 24 b^2\] What we want to do here to factor a quadratic by grouping is to multiply the coefficient of the highest order term (1) by the coefficient of the lowest order term (-24). 1*-24 = -24. Now, we factor that quantity, and look for a pair of factors which add to the coefficient of the middle term (2). -24 = -1*24 -24 = -2*12 -24 = -3*8 -24 = -4*6 -24 = -6*4 -24 = -8*3 -24 = -12*2 -24 = -24*1 Do you see any pairs of factors there which add to 2?